# Can the 3 neutrino masses really be found using SN 1987A data?

### v1

@@ Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remrkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, a theoretical linkage to possible CPT violation, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumotions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any masses $1eV/c^{2}$ would be inconsistent with upper limits from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A is too great for the method to work.
Keywords: SN 1987A, Neutrino, supernova, CPT violation, tachyon, KATRIN
@@ I Introduction
About $10^{58}$ neutrinos were emitted from the dying star Sanduleak 69 202 when it became the first supernova seen in 1987 (hence the name SN 1987A). Of those $10^{58}$ neutrinos only a couple of dozen were captured by Earth’s detectors. Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in our galaxy since neutrino detectors first existed. The early articles on the neutrino mass obtained from SN 1987A data mostly, with one exception,(Cowsik1988, ) set upper limits on its value. (Arnett1987, ; Burrows1988, ) The upper limits restriction is based on the almost universal belief that the differences in the three neutrino mass states are simply too small to use the data to find the individual masses themselves.
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) an extensive discussion of CPT violation, (2) a much improved statistical analysis, (3) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A.
@@ I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
Many theorists have regarded quantum fields giving rise to physical tachyons as being either impossible or unlikely in light of the many theoretical difficulties they create. However, unobservable $m^{2}<0$ particles seem to be OK with theorists. Thus, for most theorists the idea that symmetry-breaking in the interaction of Higgs fields with massless gauge fields leads to the production of unobservable tachyons is not objectionable. The difficulties with “physical” tachyons having an $m^{2}<0$ that can be observed include their energy spectrum being unbounded from below, their frame-dependent and unstable vacuum state, and their noncovariant commutation rules. Nevertheless, there are also theorists who are accepting of physical tachyons, and recently Paczos et al. (Paczos2024, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues. Charles Schwartz is another theorist who has challenged all the reasons why tachyons are considered unphysical. (Schwartz2022, ).
I.2 The meaning of negative searches for tachyons
@@
The question of whether $m^{2}<0$ and $v>c$ observable particles really exist is not theoretical, but strictly empirical. All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect. (Ehrlich2022, ) Most physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category. A more reasonable conclusion is that as long as not a single neutrino has ever been measured to have either $m^{2}>0$ or $v<c,$ the tachyonic neutrino hypothesis remains viable. A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA. However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is very close to zero, so we cannot tell if $m^{2}>0$ or $m^{2}<0$. Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal. Negative results for tachyonic neutrinos are not unlike negative results for searches for extraterestrials (SETI). In both cases, an appropriate response to a negative search would be “keep looking using more sensitive techniques.” Surprisinfly, however, if you ask an AI whether negative searches for tachyons make it more likely they don’t exist and then ask the same question about ET’s existence, you are illogically told yes for the former question and no for the latter.
I.3 The KATRIN experiment
@@
The KATRIN experiment seeks to find the “effective mass” of the electron antineutrino $m^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}.$ Its best value as of 2025 is negative: $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the size of the uncertainty. Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value now $0.108eV^{2}$ will be reduced. Whatever the final KATRIN results, recall that the experiment measures the effective mass of the electron neutrino and not the three individual masses comprising it, which is the primary interest of this paper.
II 3 Neutrino masses and SN 1987A
It is widely believed that we can set an upper limit on the neutrino masses based on oscillation experiments. These experiments yield values for separations in the mass squared of pairs of mass states: $\Delta m^{2}_{j,k}=m_{j}^{2}-m_{k}^{2},$ the two well-established ones being the atmospheric and solar $m^{2}$ splittings $\Delta m^{2}_{atm}$ and $\Delta m^{2}_{sol}.$
II.1 A true upper limit on Neutrino masses?
The three masses $m^{2}_{k}$ values are normally assumed to be no greater than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}\approx 0.0025eV^{2}.$ It is important to recognize this upper limit is based on a questionable assumption, namely the non-existence of any third observed oscillation frequency. In other words, since the two oscillation frequencies found for (1) $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ and (2) $\Delta m^{2}_{atm}$ are so close, they are empirically indistinguishable. If a third observed oscillation frequency did exist (perhaps the result of a fourth sterile neutrino) then the assumption of the upper limit $m_{k}^{2}<0.0025eV^{2}$ would be false. In that case, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance. Moreover, the best place to measure the three masses based on travel time might be a galactic supernova, provided the neutrino emissions from it are nearly simultaneous. No other source of a neutrino burst, even that produced by an atomic bomb, could come close to the precision made possible by the enormous travel distance to a supernova.
II.2 Supernova SN 1987A neutrinos
@@
Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987. Those detectors (Kamiokande II, IMB, and Baksan) recorded bursts of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos. The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds. This lack of synch is partially removed in most analyses by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
@@
so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy. Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened. This late light arrival, of course does not imply that neutrinos are superlumnial. Rather, the three-hour gap between the neutrino burst and the later arrival of visible light, aligns with the small interaction between neutrinos with matter. Thus, photons but not neutrinos are delayed by interactions with the supenova ejecta.
II.3 The 5 hour early LSD neutrino burst
@@
In addition to the three detectors, there was a fourth small detector known as LSD ln 1987 which was located under Mont Blanc. LSD reported a burst of five neutrinos which occurred 484 min before the other three detectors’ bursts. This early burst is therefore considered by most physicists to be a background fluctuation unrelated to SN 1987A because of its five hour early arrival. If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones ones as well. (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists.
@@
The five neutrinos observed during a 7 second time interval in LSD would if spurious represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months. In fact, the probability of a chance fluctuation giving rise to the LSD signal is extremely low, less than $1.4\times 10^{–6}.$ (Galeotti2016, ) Many observers also reject the LSD burst because no such early burst was seen in the three other detectors. However, the LSD neutrinos were all very low energy, and the other detectors could not have seen an LSD-like burst at that early time because of their higher energy thresholds. Thus, for example, the largest detector, Kamiokande II, had an energy threshold of 7 MeV, well above the energies of four of the five LSD neutrinos constituting the burst Another suspicious fact is that the main t=0 burst at 7:35 UT was not seen in LSD. However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst). Two final major problem with the LSD burst being associated with SN 1987A are (1) why the very low energies of all the 5 neutrinos comprising it ?, and (2 )why the virtual monochromaticity of the burst? In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five above objections to it are answerable, and we address the monochromaticity problem later.
III Finding $m^{2}$ of SN 1987A neutrinos
@@
The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close to simultaneously. Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of $\emph{i}ndividual$ neutrinos. In what follows $T=168,000$ years is the time photons required to reach us from SN 1987A.
@@
If a photon’s speed en route to Earth is $c=D/T$ then a neutrino’s speed would be $v=D/(t+T)\approx c(1-t/T).$ Here t is the time delay of a neutrino relative to a photon. But since there can also be such a delay if a neutrino is emitted later than other neutrinos, so the definition of t only makes sense if almost all neutrinos were emitted during a very small time window. For the neutrino speed we have v/c=1t/T.v/c=1-t/T. and also $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}.$ Setting the two versions of $v/c$ equal
@@
Eq. 3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the neutrinos will be found to lie on or near a straight line through the origin of slope M given by
@@
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line having negative slope.
@@
To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin. Any neutrinos that do not do so must either be a background event unrelated to SN 1987A, or else the result of non-simultaneous emissions. It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the neutrino masses are too close to distinguish from one another.
Before we examine what the data show, there is a philosophical issue to mention. Given that in the standard view neutrino masses are less than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ – a value so small that any spread in neutrino travel times could not possibly be large enough to observe the neutrino mass, we have two choices on how to proceed. The first is not to even bother to see what the data suggest (the ”Cremonini option”) given that we know any positive result must be an anomaly, given the small upper limt on the neutrino mass . However, we have made the second choice, namely not to make any assumptions on the size of the neutrino masses and be guided by what the data themselves show.
IV Fitting 30 data points to 3 lines

### v2

@@ Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remarkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumptions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any neutrino masses $m_{k}>1eV/c^{2}$ cannot be reconciled with upper limits on the “effective mass” from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A data is too great for the method to work. A particularly interesting piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
Keywords: SN 1987A, Neutrino, supernova, violation, tachyon, KATRIN, sterile neutrino
@@ I Introduction
About $10^{58}$ neutrinos were emitted from the dying star Sanduleak 69 202 when it became the first supernova seen in 1987 (hence the name SN 1987A). Of those $10^{58}$ neutrinos only a couple of dozen were captured by Earth’s detectors. Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in our galaxy since neutrino detectors first existed. The earliest articles on the neutrino mass obtained from SN 1987A data, with one exception,(Cowsik1988, ) set only upper limits on its value. (Arnett1987, ; Burrows1988, ; Fang1987, ). The upper limits restriction is based on the almost universal belief that the values of the three neutrino mass states are simply too small to use the SN 1987A data to find them.
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data. (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) a clue from the KATRIN experiment’s reported failure to observe sterile neutrinos. This absence surprisingly strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim. (KATRIN2025, )
@@ I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
Many theorists have regarded quantum fields giving rise to physical tachyons as being either impossible or unlikely in light of the many theoretical difficulties they create. However, unobservable $m^{2}<0$ particles seem to be OK with theorists. Thus, for most theorists the idea that symmetry-breaking in the interaction of Higgs fields with massless gauge fields leads to the production of unobservable tachyons is not objectionable. The difficulties with “physical” tachyons having an $m^{2}<0$ that can be observed include their energy spectrum being unbounded from below, their frame-dependent and unstable vacuum state, and their noncovariant commutation rules. Nevertheless, there are also theorists who are accepting of physical tachyons, and recently Paczos et al. (Paczos2024, ; Paczos2025, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues. Charles Schwartz is another theorist who has successfully challenged all the reasons why tachyons are considered unphysical, and shown how tachyons can account for many cosmological observations. (Schwartz2022, ; Schwartz2023, ).
I.2 Do negative searches make tachyons less likely?
@@
The question of whether $m^{2}<0$ and $v>c$ observable particles exist is not theoretical, but strictly empirical. All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect apart from one-time sightings. (Ehrlich2022, ) Most physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category. A more reasonable conclusion is that as long as not a single neutrino has ever been measured to have either $m^{2}>0$ or $v<c,$ the tachyonic neutrino hypothesis remains viable. A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA. However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is so close to zero, that we cannot tell if $m^{2}>0$ or $m^{2}<0$. You might think a negative hypothetical result $m^{2}=0.0\pm 0.1eV^{2}/c^{4}$ makes tachyonic neutrinos more unlikely than a previous result $m^{2}=0.0\pm 10.0eV^{2}/c^{4}$ because the allowed negative range in $m^{2}$ has been reduced a hundredfold. However, a Bayesian would realize that conclusion is false, because the allowed positive range has also been reduced by the same hundredfold amount. Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal.
I.3 KATRIN and circumventing upper limits on $m_{\nu}.$
@@
The KATRIN experiment seeks to find the “effective mass” of the electron antineutrino $m^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}.$ Its best value as of 2025 is negative: $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the size of the uncertainty. Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value will be reduced. Whatever the final KATRIN results, recall that the experiment measures the effective mass of the electron neutrino and not the three individual masses comprising it, which is the primary interest of this paper. Later we shall see how upper limits on the three neutrino masses from KATRIN and cosmological data can be circumvented.
Without any specific a priori upper limits, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance. The best baseline to measure the three masses based on their travel time is almost certainly a galactic supernova, provided we are confident the neutrino emissions from it are nearly simultaneous. An extragalactic one might seem even better were it not for the inverse square law and the very few neutrinos that could reach us. The phrase “neutrino burst” is used throughout this paper, referring to a sudden, intense emission of a vast number of neutrinos from a powerful cosmic event, most notably a core-collapse supernova lasting seconds. However we also use the phrase to refer to a seconds-long grouping of many more neutrinos than normally seen in a neutrino detector in such a time interval, which are believed to have originated from a supernova.
I.4 Supernova SN 1987A neutrinos
@@
Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987. Those detectors (Kamiokande II, IMB, and Baksan) recorded bursts of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos. The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds. This lack of synchronization is usually accounted for by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
@@
so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy. Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened. This late light arrival, of course does not imply that neutrinos are superluminal. Rather, the three-hour gap between the neutrino burst and the later arrival of visible light, aligns with the small interaction between neutrinos with matter. Thus, photons but not neutrinos are delayed by interactions with the supernova ejecta.
I.5 The 5 hour early LSD neutrino burst
@@
In addition to the three detectors, there was a fourth small detector known as LSD ln 1987 which was located under Mont Blanc. This detector was in fact better shielded from background radiation than the other three, and had a lower energy threshold, so in these two respects was more sensitive. The LSD detector reported a burst of five neutrinos which occurred 484 min before the other three detectors’ received bursts. This early burst is therefore considered by most physicists to be a background fluctuation unrelated to SN 1987A because of its five hour early arrival. If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones as well. (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists. On the other hand, there were four other reasons to be suspicious about the LSD burst besides the 5 hr early arrival, starting with it possibly being a statistical fluctuation of background.
@@
The five neutrinos observed during a 7 second time interval in LSD would, if not connected to SN1987A, represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months. In fact, the probability of a chance fluctuation giving rise to the LSD signal is extremely low, less than $1.4\times 10^{–6}.$ (Galeotti2016, ) Many observers also reject the LSD burst because no such early burst was seen in the three other detectors. However, the LSD neutrinos were all very low energy, and the other detectors could not have seen an LSD-like burst at that early time because of their higher energy thresholds. Thus, for example, the largest detector, Kamiokande II, had an energy threshold of 7 MeV, well above the energies of four of the five LSD neutrinos constituting the burst Another suspicious fact is that the main t=0 burst at 7:35 UT was not seen in LSD. However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst). Two final major problem with the LSD burst being associated with SN 1987A are (1) the very low energies of all the 5 neutrinos comprising it, and (2 ) their nearly equal energies or the virtual monochromaticity of the burst. In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five objections to it are answerable, but we address the monochromaticity problem later.
II Finding $m^{2}$ of SN 1987A neutrinos
@@
The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close enough to simultaneously. Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of the neutrinos. In what follows $T=168,000$ years is the time photons required to reach us from SN 1987A.
@@
If a photon’s speed en route to Earth is $c=D/T$ then a neutrino’s speed would be $v=D/(t+T)\approx c(1-t/T).$ Here t is the time delay of a neutrino relative to a photon. But since there can also be such a delay if a neutrino is emitted later than other neutrinos, so the definition of t only makes sense if almost all neutrinos were emitted during a very small time window. For the neutrino speed we have v/c=1t/T.v/c=1-t/T. and also as an excellent approximation $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}$. Setting the two versions of $v/c$ equal
@@
Eq. 3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the data point for each neutrino will be found to lie on or near a straight line through the origin of slope M given by
@@
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin having negative slope.
@@
To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) non-sterile neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin. Any neutrinos that do not do so must either be a background event unrelated to SN 1987A, or else the result of non-simultaneous emissions. It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the lines (or neutrino masses) are too close to distinguish from one another.
III Fitting 30 data points to 3 lines